# Efficient tomography of microwave photonic cluster states

### v2

@@ I INTRODUCTION
To fully characterize an experimentally-generated cluster state, its many-body density matrix needs to be reconstructed from a complete set of measurements. The reconstructed density matrix can then be used to calculate such metrics as the localizable entanglement [24, 25], which can evaluate the usefulness of an entangled state for measurement-based quantum computation [26]. However, reconstructing the density matrix of a large entangled state has remained challenging because the conventional quantum state tomography becomes exponentially costly as the number of qubits increases. To circumvent this difficulty, large-scale cluster states have been characterized in one of two ways: evaluating an entanglement witness [15, 17, 18] or performing a quantum process tomography of each operation in the entanglement generation protocol [16, 19]. In the first approach, theoretical work is required to construct a witness operator for each purpose, such as certifying genuine multipartite entanglement [27, 28] and lower-bounding the quantum state fidelity [18] or localizable entanglement [29]. The second approach evaluates the individual steps of the protocol in isolation, which means that a numerical extrapolation is required to evaluate the entire protocol. Recently, Ref. 23 demonstrated a tomography of a 20-qubit quasi-two-dimensional microwave photonic cluster state by implementing an iterative maximum-likelihood algorithm using a matrix-product-operator (MPO) representation of the density matrix [30]. However, an algorithm with a provably fast convergence and a way to propagate the statistical uncertainties of the data is more desirable.
@@
Here, we propose and experimentally demonstrate an efficient tomography method for reconstructing the many-body density matrix of a sequentially emitted chain of entangled photonic qubits. This is made possible by the fact that the exponentially large density matrix can be parameterized efficiently using an MPO representation even if there are coherent and incoherent errors in the photon emission process. To demonstrate the tomography method, we use a superconducting qubit to generate linear cluster states of up to 35 microwave photonic qubits in the discrete-variable basis. We then use a quantum-limited amplifier to measure the quadrature observables of the photonic qubits and efficiently obtain the complete set of correlations among five consecutive qubits. The correlation data is first used to verify that the density matrix of the generated photons can be represented by an MPO with a bond dimension of four. Then, we use a gradient-based least-squares algorithm to efficiently fit the MPO to the measured correlations and propagate the measurement uncertainties to the parameters of the MPO. Using the reconstructed MPO representations of the cluster states, we calculate the quantum state fidelity and localizable entanglement, which are found to be significantly larger than the lower bounds determined by stabilizer measurements [28, 29]. We demonstrate that the entanglement in the 10-qubit cluster state persists for up to a length of seven consecutive photons. We also find that this length decreases for the larger cluster states, which contradicts the numerical extrapolation of the results for the smaller cluster states. Our observation suggests that the successful generation of a smaller cluster state does not guarantee the generation of a larger cluster state, highlighting the importance of efficient tomography methods for large entangled states.
@@ II EFFICIENT TOMOGRAPHY
Two factors contribute to the exponential difficulty of performing a quantum state tomography of an entangled state of a large number of qubits. The first is the large number of parameters in a multi-qubit density matrix, which increases exponentially with the number of qubits. The second is the large number of possible measurement settings, i.e., the choice of measurement basis for each qubit, which also increases exponentially. In the first part of this section, we resolve the first difficulty by showing that a sequentially generated entangled state can be represented by an MPO whose number of parameters only grows linearly with the number of qubits. Then, we show that the second difficulty can be resolved using the fact that an MPO can be reconstructed from local correlation measurements, which only require a constant number of measurement settings regardless of the number of qubits.
@@ II.1 Efficient parameterization of sequentially generated entanglement
where $\rho_{i_{1},\ldots,i_{N}}$ is a $4\times\cdots\times 4$ real tensor, and $\hat{P}^{(0)}_{s}=\hat{I}_{s}$, $\hat{P}^{(1)}_{s}=\hat{X}_{s}$, $\hat{P}^{(2)}_{s}=\hat{Y}_{s}$, and $\hat{P}^{(3)}_{s}=\hat{Z}_{s}$ are the Pauli operators on the $s$th qubit for $s=1,\ldots,N$. In the conventional quantum state tomography, each parameter $\rho_{i_{1},\ldots,i_{N}}$ is estimated from an $N$-qubit correlation measurement as $\rho_{i_{1},\ldots,i_{N}}=\langle\hat{P}^{(i_{1})}_{1}\cdots\hat{P}^{(i_{N})}_{N}\rangle$. Taking into account the unit-trace constraint, $\rho_{0,\ldots,0}=\operatorname{Tr}[\hat{\rho}]=1$, the number of free parameters is $14^{N}-$ 1, which makes this procedure exponentially difficult as the number of qubits increases.
In the MPO representation, the $4\times\cdots\times 4$ tensor $\rho_{i_{1},\ldots,i_{N}}$ is decomposed into a contraction of $N$ tensors: a $4\times D$ tensor $\mathsf{A}_{1}$, $D\times 4\times D$ tensors $\mathsf{A}_{2},\ldots,\mathsf{A}_{N-1}$, and a $D\times 4$ tensor $\mathsf{A}_{N}$ [31]. The contraction can be expressed as matrix multiplications by slicing each tensor along the 4-dimensional axis:
@@
The MPO representation has been used to efficiently parameterize and characterize a quantum state generated by a system with local interactions such as a linear ion trap [34, 35]. Here, we show that a sequentially-emitted chain of entangled photonic qubits can also be efficiently represented by an MPO because of the locality in the time dimension. We expand on Ref. 36, which showed that a chain of photonic qubits generated by a $d$-level system (qudit) can be represented by a matrix product state (MPS) with a bond dimension of at most $d$ if the qudit–photon system remains in a pure state. We use the concepts of mixed-state quantum circuits [37] and tensor networks [38] to confirm the suggestion in Ref. 36 that an MPO can be used to allow for mixed states and decoherence.
@@
Figure 1: Efficient tomography of sequentially-emitted entangled photonic qubits. (a) Sequential emission of entangled photonic qubits by a $d$-level system (qudit). (b) Mixed-state quantum-circuit representation, which can be interpreted as a tensor network corresponding to the density matrix of the emitted photonic qubits. (c) Matrix product operator (MPO) obtained by contracting the tensors in groups. This is an efficient parameterization of the many-body density matrix of the generated photonic qubits because the number of parameters only increases proportionally with the number of photonic qubits, $N$. (d) Efficient choice of measurement bases which makes the number of measurement settings independent of $N$. Here, we measure the complete set of local correlations of up to five consecutive photonic qubits, which is sufficient for reconstructing the density matrix of a linear cluster state.
[…]
@@
This quantum circuit can be interpreted as a tensor-network diagram corresponding to the density matrix of the emitted photonic qubits. In a tensor network diagram, a node with $n$ edges represents a tensor with $n$ indices, an edge connecting two nodes represents a tensor contraction, and a dangling edge represents an index of the resulting tensor. Then, the tensors can be contracted in groups as shown by the dashed orange boxes in Fig. 1(b) to obtain a tensor network shown in Fig. 1(c), which corresponds to Eq. (2) and represents an MPO. The dimension of the tensor contractions represented by the horizontal edges corresponds to the bond dimension of the MPO. Therefore, the $N$-qubit density matrix of the generated photonic qubits can be efficiently parameterized using an MPO with a bond dimension of at most $D=d^{2}$, whose number of parameters only grows linearly with $N$. Notably, this representation is efficient even if additional energy levels of the photon emitter are unintentionally involved, under the reasonable assumption that the number of the participating energy levels does not increase exponentially with $N$. Thus, the MPO representation allows one to perform the quantum state tomography of a sequentially generated entanglement without estimating each of the exponentially many parameters of the density matrix.
@@ II.2 Efficient choice of measurement settings
Our strategy for choosing an efficient set of correlation measurements is based on Ref. 34, which showed that an $N$-qubit state represented by an MPO can be reconstructed from its local reductions if the “reconstructibility condition” described below is satisfied. A local reduction of a chain of qubits is defined as the state obtained by selecting $L$ consecutive qubits and ignoring the others. Because the local reduction is an $L$-qubit state, it is determined by the complete set of $L$-qubit correlation measurements. Taking $L=5$ as an example, the local correlations can be denoted as
@@
where $a,b,c,d,e\in\{0,1,2,3\}$ and $s=1,\ldots,N-4$. The MPO representation of the $N$-qubit state can be reconstructed from the set of local correlations by following the procedure introduced in Ref. 34. Appendix G gives an explicit formula for the reconstructed MPO in terms of $C^{(a,b,c,d,e)}_{s}$. The reconstructibility condition is fulfilled by the vast majority of MPOs whose bond dimension $D$ satisfies
@@
[…]
@@ III.1 Generating a microwave photonic cluster state
Figure 2: Generation and measurement of a microwave photonic linear cluster state. (a) Photograph of the disassembled (front) and assembled (back) microwave photon sources, which consist of a superconducting transmon qubit and a microwave resonator on a silicon substrate inside an aluminum cylindrical cavity. The threaded connector with a long center pin is used to couple a coaxial cable to the resonator. (b) Circuit model of the photon source and a schematic of the measurement chain, which includes a Josephson parametric amplifier (JPA) for phase-sensitive quantum-limited amplification. (c) Energy-level diagram of the photon source. The numbered arrows correspond to the steps comprising the conditional photon emission operation. (d) Measured mode function $u(t)$ of the photonic qubit. (e) Probability density $p(q_{1},q_{2})$ of measuring the quadrature values $q_{1}$ and $q_{2}$ for the first and second photonic qubits, respectively, of the generated five-qubit cluster state. The solid and dashed lines are the contour lines for the measured data and an ideal cluster state under the effect of the same measurement inefficiency, respectively. (f) Pulse sequence for generating an $N$-qubit linear cluster state. (g) Measured photon flux of the cluster states. The traces are offset vertically by units of 20 $\mu$s-1 for clarity.
To generate a linear cluster state, we use a protocol based on the proposal in Ref. 12, which uses an electron confined in a semiconductor quantum dot to generate a linear cluster state of optical photons. This protocol inspired the experimental demonstrations using a dark exciton [16], an InGaAs quantum dot [17], a Rubidium atom [18], and superconducting qubits [19, 22]. Our protocol is based on a microwave-activated photon–qudit interaction and is different from the previous works in the microwave domain in the sense that it does not use any frequency-tunable qubit or tunable coupler. This protocol is promising for increasing the fidelity and scale of the cluster states because tuning mechanisms tend to increase the decoherence rates and wiring complexity of the photon source.
@@
Figure 2(f) shows the pulse sequence for generating an $N$-qubit linear cluster state (see Appendix D for more details). The transmon is initialized to the ground state using a quantum nondemolition readout and post-selection. Then, we repeat the 200-ns cycle consisting of a 20-ns $\pi/2$ rotation in the qubit subspace $\{\lvert g0\rangle,\lvert e0\rangle\}$ and a photon emission conditioned on the qubit being in $\lvert e0\rangle$. The conditional photon emission is composed of two 20-ns $\pi$-pulses which induce $\lvert e0\rangle$$\lvert f0\rangle$ and $\lvert g0\rangle$$\lvert e0\rangle$ transitions in this order, followed by a $\lvert f0\rangle$$\lvert g1\rangle$ drive, which is a chirped raised-cosine pulse [46] with a duration of 100 ns. After repeating the 200-ns cycle $N-1$ times, another $\pi/2$ rotation is performed and the transmon is disentangled from the generated photons by completely emitting its excitation as a photon. We used this pulse sequence to generate $N$-qubit cluster states for $N=5$, 10, 15, 20, 25, 30, and 35. Figure 2(g) shows the measured photon flux of the cluster states.
@@ III.2 Measuring the microwave photonic qubits
We measure the quadrature observables of the microwave photonic qubits using a flux-driven Josephson parametric amplifier (JPA) [49]. By driving the JPA with a 140-ns pulse at twice the signal frequency, the signal can be phase-sensitively amplified. For each qubit, either the $\hat{q}_{s}$ or $\hat{p}_{s}$ quadrature is amplified by switching the phase of the pump pulse between 0 and $\pi$. The quadrature values are obtained by calculating the overlap integral between the amplified waveform and the mode function $u(t)$ of the photonic qubit. Figure 2(e) shows the two-dimensional histogram of the quadrature values measured for the first and second photonic qubits of the five-qubit cluster state. The visible correlation between $\hat{q}_{1}$ and $\hat{q}_{2}^{2}$ corresponds to the Pauli-basis correlation $\langle\hat{X}_{1}\hat{Z}_{2}\rangle=1$, which is characteristic of a cluster state. Whereas this histogram is not corrected for the measurement inefficiency, we correct the data used in the following analysis using the estimated measurement efficiency of $0.391\pm 0.004$ (see Appendix E for details). We also numerically correct for the phase offsets between the generated photonic qubits and the measured quadratures by maximizing the expectation values of the stabilizer operators of the cluster state (see Appendix F for details).
To measure a five-qubit correlation, we choose for each photonic qubit either $\hat{q}_{s}$ or $\hat{p}_{s}$ as the measurement basis, perform $10^{7}$ measurements with a repetition period of 10 $\mu$s, and calculate the set of moments
@@
for a pulse mode containing up to one photon. Only $322^{5}=$ 32 measurement settings, i.e., the combination of the JPA phase for each photonic qubit, are required to measure the complete set of five-qubit local correlations for the entire photon chain.

### v3

@@ I INTRODUCTION
To fully characterize an experimentally-generated cluster state, its many-body density matrix needs to be reconstructed from a complete set of measurements. The reconstructed density matrix can then be used to calculate such metrics as the localizable entanglement [64, 47], which can evaluate the usefulness of an entangled state for measurement-based quantum computation [59]. However, reconstructing the density matrix of a large entangled state has remained challenging because the conventional quantum state tomography becomes exponentially costly as the number of qubits increases. To circumvent this difficulty, large-scale cluster states have been characterized in one of two ways: evaluating an entanglement witness [68, 25, 57] or performing a quantum process tomography of each operation in the entanglement generation protocol [53, 13]. In the first approach, theoretical work is required to construct a witness operator for each purpose, such as certifying genuine multipartite entanglement [60, 58] and lower-bounding the quantum state fidelity [57] or localizable entanglement [43]. The second approach evaluates the individual steps of the protocol in isolation, which means that a numerical extrapolation is required to evaluate the entire protocol. This approach also requires the assumption that the photon emission process executed during the protocol is identical to when it is executed in isolation (Table 1). Recently, Ref. 44 demonstrated a tomography of a 20-qubit quasi-two-dimensional microwave photonic cluster state by implementing an iterative maximum-likelihood algorithm using a matrix-product-operator (MPO) representation of the density matrix [7]. However, for a large problem, the convergence of the iterative maximum-likelihood algorithm is significantly slower than that of a direct optimization algorithm [41]. Moreover, the convergence cannot be guaranteed without further reducing the convergence rate [50].
Table 1: Comparison of tomography methods for entangled photonic qubits.
Conventional tomography Process tomography and numerical extrapolation [53, 13] MPO reconstruction
Number of parameters to estimate $4^{N}-1$ $O(1)$ $O(N)$
Assumes sequentially emitted photons? No Yes Yes
Assumes identical emission process? No Yes No
Table 2: Comparison of tomography methods based on MPO reconstruction.
Direct inversion [8] Iterative maximum likelihood [7, 44] This work
Requires iterative optimization of MPO? No Yes (slow) Yes (fast)
Robust against statistical noise of high-order correlations? No No Yes
Guarantees positivity of reconstructed MPO? No Yes No
@@
Here, we propose and experimentally demonstrate an efficient tomography method for reconstructing the many-body density matrix of a sequentially emitted chain of entangled photonic qubits. This is made possible by the fact that the exponentially large density matrix can be parameterized efficiently using an MPO representation even if there are coherent and incoherent errors in the photon emission process. To demonstrate the tomography method, we use a superconducting qubit to generate linear cluster states of up to 35 microwave photonic qubits in the discrete-variable basis. We then use a quantum-limited amplifier to measure the quadrature observables of the photonic qubits and efficiently obtain the complete set of correlations among five consecutive qubits. An MPO is then fitted to the measured correlations using the Gauss–Newton weighted least-squares algorithm, which is known to converge linearly or better given a good initial guess. The weighted fitting makes our method robust against the large statistical noise inherent in high-order correlation measurements and enables us to propagate the measurement uncertainties to the parameters of the MPO. To obtain the good initial guess, we derive an approximate reconstruction formula for the MPO in terms of the measured correlations by reformulating the direct inversion procedure introduced in Ref. 8. This reformulation also provides a method to directly estimate the required bond dimension of the MPO representation from the measured correlations. These features make our tomography method an attractive alternative to methods based on direct inversion [8] or the iterative maximum-likelihood algorithm [7, 44]. Table 2 summarizes the differences between these tomography methods.
Using the reconstructed MPO representations of the cluster states, we calculate the quantum state fidelity and localizable entanglement. The localizable entanglement is found to be significantly larger than the lower bound determined by stabilizer measurements [43]. We demonstrate that the entanglement in the 10-qubit cluster state persists for up to a length of seven consecutive photons. We also find that this length decreases for the larger cluster states, which contradicts the numerical extrapolation of the results for the smaller cluster states. Our observation suggests that the successful generation of a smaller cluster state does not guarantee the generation of a larger cluster state, highlighting the importance of efficient tomography methods for large entangled states.
@@ II EFFICIENT TOMOGRAPHY
In this section, we first show that the process tensor representing the sequential emission of photonic qubits can be transformed into an MPO. This means that the many-body density matrix of the emitted photonic qubits can be efficiently represented using a linear number of parameters with respect to the number of qubits. However, a parameter in this representation does not directly correspond to a measurable quantity of the photon chain. To realize an efficient tomography, one needs to choose a small set of measurable quantities to which these parameters can be fitted. In the second part of this section, we describe how a linear cluster state can be tomographically reconstructed from a set of local correlations, which can be measured in parallel using a constant number of measurement settings regardless of the number of qubits.
@@ II.1 Efficient parameterization of sequentially generated entanglement
where $\rho_{i_{1},\ldots,i_{N}}$ is a $4\times\cdots\times 4$ real tensor, and $\hat{P}^{(0)}_{s}=\hat{I}_{s}$, $\hat{P}^{(1)}_{s}=\hat{X}_{s}$, $\hat{P}^{(2)}_{s}=\hat{Y}_{s}$, and $\hat{P}^{(3)}_{s}=\hat{Z}_{s}$ are the Pauli operators on the $s$th qubit for $s=1,\ldots,N$. In the conventional quantum state tomography, each parameter $\rho_{i_{1},\ldots,i_{N}}$ is estimated from an $N$-qubit correlation measurement as $\rho_{i_{1},\ldots,i_{N}}=\langle\hat{P}^{(i_{1})}_{1}\cdots\hat{P}^{(i_{N})}_{N}\rangle$. Taking into account the unit-trace constraint, $\rho_{0,\ldots,0}=\operatorname{Tr}[\hat{\rho}]=1$, the number of free parameters is $4^{N}-1$, which makes this procedure exponentially difficult as the number of qubits increases.
In the MPO representation, the $4\times\cdots\times 4$ tensor $\rho_{i_{1},\ldots,i_{N}}$ is decomposed into a contraction of $N$ tensors: a $4\times D$ tensor $\mathsf{A}_{1}$, $D\times 4\times D$ tensors $\mathsf{A}_{2},\ldots,\mathsf{A}_{N-1}$, and a $D\times 4$ tensor $\mathsf{A}_{N}$ [17]. The contraction can be expressed as matrix multiplications by slicing each tensor along the four-dimensional axis:
@@
The MPO representation has been used to efficiently parameterize and characterize a quantum state generated by a system with local interactions such as a linear ion trap [8, 33]. Here, we show that a sequentially-emitted chain of entangled photonic qubits can also be efficiently represented by an MPO because of the locality in the time dimension. We expand on Ref. 52, which showed that a chain of photonic qubits generated by a $d$-level system (qudit) can be represented by a matrix product state (MPS) with a bond dimension of at most $d$ if the qudit–photon system remains in a pure state. We use the concepts of mixed-state quantum circuits [1] and tensor networks [14] to confirm the suggestion in Ref. 52 that an MPO can be used to allow for mixed states and decoherence.
@@
Figure 1: Efficient tomography of sequentially-emitted entangled photonic qubits. (a) Sequential emission of entangled photonic qubits by a $d$-level system (qudit). (b) Mixed-state quantum-circuit representation, which can be interpreted as a tensor network corresponding to the density matrix of the emitted photonic qubits. (c) Matrix product operator (MPO) obtained by contracting the tensors in groups. This is an efficient parameterization of the many-body density matrix of the generated photonic qubits because the number of parameters only increases proportionally with the number of photonic qubits, NN. (d) Efficient choice of measurement bases which makes the number of measurement settings independent of $N$. Here, we measure the complete set of local correlations of up to five consecutive photonic qubits, which is sufficient for reconstructing the density matrix of a linear cluster state.
Figure 1(a) depicts a qudit sequentially emitting a chain of photonic qubits. This process can be modeled using the mixed-state quantum circuit shown in Fig. 1(b). The initial density matrix of the qudit is represented by a $d^{2}$-dimensional real vector $\rho_{0}$ by assuming a Hermitian matrix basis such as the generalized Gell-Mann matrices [26]. Then, the effect of a gate operation on the qudit, including any coherent and incoherent errors, is to multiply the vector by a $d^{2}\times d^{2}$ process matrix $\mathcal{G}_{s}$. Similarly, a conditional photon emission from the qudit is a gate operation between the qudit and the photonic qubit and is represented by a $d^{2}\times 2^{2}\times d^{2}\times 2^{2}$ process tensor $\mathcal{E}_{s}$. Note that the initial state of the photonic qubit is chosen to be the vacuum state $\lvert 0\rangle\!\langle 0\rvert$ without loss of generality. A chain of $N$ photonic qubits is generated by alternating between the qudit gates $\mathcal{G}_{s}$ and photon emissions $\mathcal{E}_{s}$, which may differ each time as a part of the experimental protocol or because of its imperfect execution. Finally, an entangled state of the $N$ photonic qubits is obtained by ignoring the qudit state at the end of the protocol, which is equivalent to taking the partial trace over the qudit.
@@
This quantum circuit can be interpreted as a tensor-network diagram corresponding to the density matrix of the emitted photonic qubits. In a tensor-network diagram, a node with $n$ edges represents a tensor with $n$ indices, an edge connecting two nodes represents a tensor contraction, and a dangling edge represents an index of the resulting tensor. Then, the tensors can be contracted in groups as shown by the dashed orange boxes in Fig. 1(b) to obtain a tensor network shown in Fig. 1(c), which corresponds to Eq. (2) and represents an MPO. The dimension of the tensor contractions represented by the horizontal edges corresponds to the bond dimension of the MPO. Therefore, the $N$-qubit density matrix of the generated photonic qubits can be efficiently parameterized using an MPO with a bond dimension of at most $D=d^{2}$, whose number of parameters only grows linearly with $N$. Notably, this representation is efficient even if additional energy levels of the photon emitter are unintentionally involved, under the reasonable assumption that the number of the participating energy levels does not increase exponentially with $N$. Thus, the MPO representation allows one to perform the quantum state tomography of a sequentially generated entanglement without estimating each of the exponentially many parameters of the density matrix.
@@ II.2 Efficient choice of measurement settings
Our strategy for choosing an efficient set of correlation measurements is based on Ref. 8, which showed that an $N$-qubit state represented by an MPO can be reconstructed from its local reductions if the “reconstructibility condition” described below is satisfied. A local reduction of a chain of qubits is defined as the state obtained by selecting $L$ consecutive qubits and ignoring the others. Because the local reduction is an $L$-qubit state, it is determined by the complete set of $L$-qubit correlation measurements. Taking $L=5$ as an example, the local correlations can be denoted as
@@
where $a,b,c,d,e\in\{0,1,2,3\}$ and $s=1,\ldots,N-4$. The MPO representation of the $N$-qubit state can be reconstructed from the set of local correlations by following the procedure introduced in Ref. 8. Appendix G gives an explicit formula for the reconstructed MPO in terms of $C^{(a,b,c,d,e)}_{s}$. The reconstructibility condition is fulfilled by the vast majority of MPOs whose bond dimension $D$ satisfies
@@
where $\lfloor\cdot\rfloor$ denotes the floor function (see Appendix G for the full statement of the reconstructibility condition). The exceptional cases form a zero-volume subset within the space of all MPOs with a bond dimension of $D$. In these cases, the MPO is unreconstructible because one or more of the linear systems of equations which appear in the reconstruction procedure has no solution.
@@ III.1 Generating a microwave photonic cluster state
[…]
To generate a linear cluster state, we use a protocol based on the proposal in Ref. 36, which uses an electron confined in a semiconductor quantum dot to generate a linear cluster state of optical photons. This protocol inspired the experimental demonstrations using a dark exciton [53], an InGaAs quantum dot [25], a rubidium atom [57], and superconducting qubits [13, 19]. Our protocol is based on a microwave-activated photon–qudit interaction and is different from the previous works in the microwave domain in the sense that it does not use any frequency-tunable qubit or tunable coupler. This protocol is promising for increasing the fidelity and scale of the cluster states because tuning mechanisms tend to increase the decoherence rates and wiring complexity of the photon source.
@@
Figure 2(f) shows the pulse sequence for generating an $N$-qubit linear cluster state (see Appendix D for more details). The transmon is initialized to the ground state using a quantum nondemolition readout and post-selection. Then, we repeat the 200-ns cycle consisting of a 20-ns $\pi/2$ rotation in the qubit subspace $\{\lvert g0\rangle,\lvert e0\rangle\}$ and a photon emission conditioned on the qubit being in $\lvert e0\rangle$. The conditional photon emission is composed of two 20-ns $\pi$-pulses which induce $\lvert e0\rangle$$\lvert f0\rangle$ and $\lvert g0\rangle$$\lvert e0\rangle$ transitions in this order, followed by a $\lvert f0\rangle$$\lvert g1\rangle$ drive, which is a chirped raised-cosine pulse [23] with a duration of 100 ns. After repeating the 200-ns cycle N1N-1 times, another $\pi/2$ rotation is performed and the transmon is disentangled from the generated photons by completely emitting its excitation as a photon. We used this pulse sequence to generate $N$-qubit cluster states for $N=5$, 10, 15, 20, 25, 30, and 35. Figure 2(g) shows the measured photon flux of the cluster states.
@@ III.2 Measuring the microwave photonic qubits
Each photonic qubit generated by this protocol is defined in the $\{\lvert 0\rangle,\lvert 1\rangle\}$ subspace of a microwave pulse mode. Such a qubit can be measured in the Pauli $\hat{Z}$ basis using a microwave photon detector [24, 30, 12, 6]. Pauli $\hat{X}$ and $\hat{Y}$ measurements can be realized by absorbing the photon into a superconducting qubit, which requires temporally modulating the coupling strength to match the mode function of the photon [31, 32]. Here, we take an experimentally simpler approach of measuring the quadrature observables $\hat{q}_{s}\coloneqq(\hat{a}_{s}^{\dagger}+\hat{a}_{s})/\sqrt{2}$ and $\hat{p}_{s}\coloneqq i(\hat{a}_{s}^{\dagger}-\hat{a}_{s})/\sqrt{2}$, where $\hat{a}_{s}$ is the annihilation operator of the pulse mode for the $s$th photonic qubit. The data analysis methods in this work are directly applicable to experiments in the optical domain because the same quantities can be measured using optical homodyne detection [37].
We measure the quadrature observables of the microwave photonic qubits using a flux-driven Josephson parametric amplifier (JPA) [30]. By driving the JPA with a 140-ns pulse at twice the signal frequency, the signal can be phase-sensitively amplified. For each qubit, either the $\hat{q}_{s}$ or $\hat{p}_{s}$ quadrature is amplified by switching the phase of the pump pulse between 0 and $\pi$. The quadrature values are obtained by calculating the overlap integral between the amplified waveform and the mode function $u(t)$ of the photonic qubit measured with the method described in Appendix E.1. Figure 2(e) shows the two-dimensional histogram of the quadrature values measured for the first and second photonic qubits of the five-qubit cluster state. The visible correlation between $\hat{q}_{1}$ and $\hat{q}_{2}^{2}$ corresponds to the Pauli-basis correlation $\langle\hat{X}_{1}\hat{Z}_{2}\rangle=1$, which is characteristic of a cluster state. Whereas this histogram is not corrected for the measurement inefficiency, we correct the data used in the following analysis using the estimated measurement efficiency of $0.391\pm 0.004$ (see Appendix E for details). We also numerically correct for the phase offsets between the generated photonic qubits and the measured quadratures by maximizing the expectation values of the stabilizer operators of the cluster state (see Appendix F for details).
@@
for a pulse mode containing up to one photon. Only $2^{5}=32$ measurement settings, i.e., the combination of the JPA phase for each photonic qubit, are required to measure the complete set of five-qubit local correlations for the entire photon chain.